angela’s aunt is two more than three times her age. The total of their ages is 54. How old is Angela and how old is her aunt?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

This one is a duplicate.
https://www.algebra.com/algebra/homework/word/age/Age_Word_Problems.faq.question.1174507.html


-------------------------------------------------
angela’s aunt is two more than three times her age. The total of their ages is 54.
--------------------------------------------------

angela=13 and aunt=41

Answer by ikleyn(52790)

.

    Angela +  aunt  = 54


     A     + (3A+2) = 54


          4A        = 54 - 2 = 52


           A                 = 52/4 = 13


ANSWER.  Angela is 13 years old.  Her aunt is  3*13 + 2 = 41 years old.


CHECK.   13 + 41 = 54.    ! Correct !

Solved.

The answer by @josgarithmetic is WRONG --- so ignore it.